Circular No. 8145

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2003 L1 (SCOTTI) J. V. Scotti, Lunar and Planetary Laboratory, University of Arizona,

reports the discovery of a comet on CCD images taken with the Spacewatch 0.9-m f/3 reflector at Kitt Peak (discovery observation below), showing a coma of diameter 6′′ and a faint tail ∼ 0′.62 long in p.a. 273^{o}. Images taken by A. S. Descour on June 5.3 UT with the 1.8-m f/2.7 Spacewatch reflector also show a tail, and June 7.2 images by Scotti with the larger instrument show the tail 0′.30 long in p.a. 273^{o}.

2003 UT  α_{2000}  δ_{2000}  m_{1}
June 4.20655  14^{h}48^{m}10.^{s}53  −15^{o}24′45′′.0  19.7

The available astrometry, the following preliminary parabolic orbital elements, and an ephemeris appear on MPEC 2003-L33. The comet is very probably of short period.

T = 2002 Jan. 11.952 TT  ω = 270.^{o}575
Ω = 226.498 }2000.0
q = 2.66888 AU  i = 12.403

COMET P/2003 K2 (CHRISTENSEN) This comet appears to be of short period, the following orbital elements

taken from MPEC 2003-L31:

T = 2003 Apr. 8.059 TT  ω = 346.^{o}655
e = 0.84062  Ω = 93.550 }2000.0
q = 0.55568 AU  i = 10.152
a = 3.48647 AU n^{o} = 0.151400 P = 6.51 years

COMET C/2003 K4 (LINEAR)

Improved parabolic orbital elements from MPEC 2003-L29:

T = 2004 Oct. 13.497 TT  ω = 198.^{o}563
Ω = 18.748 }2000.0
q = 1.02128 AU  i = 134.279
2003 June 7  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green